# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ 5.1. Main estimates
(5.15) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-{t\wedge\tau_{R}})}\right]+\mathfrak{c}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r\\ =&\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-{t\wedge\tau_{R}})}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-\mathbb{E}\left[e^{\gamma(s-(t\wedge\tau_{R}))}\right]\right).\end{split}$
@@
(5.16) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}e^{\gamma(r-t)}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-t)}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-e^{\gamma(s-t)}\right).\end{split}$
By [14, Lemma 4.4], there exists a subsequence $\{n_{k}\}$ such that $Q_{n_{k}}X^{(n_{k}),x_{n_{k}}}\rightharpoonup^{\ast}X^{x}$ weakly in $L^{\infty}([0,T];L^{2}(\Omega;H))$ and weakly in $L^{\alpha}([0,T]\times\Omega,{\mathcal{B}}{\mathcal{F}},\mathrm{d}t\otimes\mathbb{P};V)$ as $k\to\infty$, where
:=${\mathcal{B}}{\mathcal{F}}:=\{A\subset[0,T]\times\Omega\,\colon\,\forall t\in[0,T],A\cap([0,t]\times\Omega)\in{\mathcal{B}}([0,t])\otimes{\mathcal{F}}_{t}\}$
is the $\sigma$-field of progressively measurable sets on $[0,T]\times\Omega$. By Lemma B.2, combined with Fatou’s lemma, we can pass to the limit $k\to\infty$ and get
(5.17) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}e^{\gamma(r-t)}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}\|_{H}^{\beta+2}e^{\gamma(s-t)}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-e^{\gamma(s-t)}\right).\end{split}$
for $\mathrm{d}t$-almost every $0\leq s\leq t\leq T$. The proof is complete. As the right-hand side depends continuously on $s$ and $t$, we get the result for every $0\leq s\leq t\leq T$ by a simple approximation argument. ∎
By investigating the above proof, we get, by similar arguments, the following:
@@
Assume that Hypotheses (A)(B)(C)(D)(E), and (F) hold true. Let $T>0$ be fixed. Let $\alpha\geq 2,\beta\geq 0$ be as in Hypothesis (C). Then for any initial datum $X_{0}\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, and every $0\leq s\leq t\leq T$, the solution $(X_{t})_{t\in[0,T]}$ to (2.1) satisfies
(5.18) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\leq\mathbb{E}\left[\|X_{s}\|_{H}^{\beta+2}\right]+\mathfrak{c}(t-s),\end{split}$
where $\delta_{1}>0$ is as in Hypothesis (B), and $\delta_{2}>0$ is as in Hypothesis (C), and where $\lambda_{0}\in(0,1]$ and $\mathfrak{c}\geq 0$ are as in Definition 5.1, where $\lambda_{0}=1$ and $\mathfrak{c}=0$ if and only if $B\equiv G\equiv 0$. In particular, $X\in L^{\infty}([0,T];L^{\beta+2}(\Omega,{\mathcal{F}},\mathbb{P};H))$, and $\mathfrak{c}$ does not depend on $T$.
@@ 5.2. Properties of the semigroup
Assume that Hypotheses (A)(B)(C)(D)(E)(F), and (G) are valid. Let $\varepsilon>0$, $F\in\operatorname{Lip}_{b}(H)$, $x\in H$. Then there exists $\delta>0$ such that for any $y\in H$ with $\|x-y\|<\delta$, and any $t\geq 0$, we have that
@@
Note that if $C_{2}=0$, the $e$-property follows as in [34, Equation (5.4)]. Let $\varepsilon>0$, $x,y\in H$, $\|x-y\|<\delta$, $F\in\operatorname{Lip}_{b}(H)$, $t\geq 0$. Let $(X^{x}_{t})_{t\geq 0}$, $(X^{y}_{t})_{t\geq 0}$ denote the solutions to (2.1) with $X_{0}^{x}=x$, and $X^{y}_{0}=y$, respectively. W.l.o.g. $\|F\|_{\infty}>0$ and $\|F\|_{\operatorname{Lip}}>0$. We claim that
(5.19) $|P_{t}F(x)-P_{t}F(y)|<\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}\left(\left\{\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H}\geq\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}\right\}\right).$
Let $J_{\varepsilon}:=\{|F(X_{t}^{x})-F(X_{t}^{y})|<\frac{\varepsilon}{3}\}$. Then,
(5.20) $\begin{split}&|P_{t}F(x)-P_{t}F(y)|\\ \leq&\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|]\\ =&\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|1_{J_{\varepsilon}}]+\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|1_{J_{\varepsilon}^{c}}]\\ \leq&\frac{\varepsilon}{3}+\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|1_{J_{\varepsilon}^{c}}]\\ \leq&\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}(J_{\varepsilon}^{c})\\ =&\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}\left(\left\{|F(X_{t}^{x})-F(X_{t}^{y})|\geq\frac{\varepsilon}{3}\right\}\right)\\ \leq&\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}\left(\left\{\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H}\geq\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}\right\}\right),\end{split}$
which proves (5.19). Let us investigate the tail probability. By Itô’s formula and Hypothesis (C), we have for $0\leq s\leq t$ and $\gamma\in(0,\delta_{2})$ as in (5.10) that
$\|X_{t}^{x}-X_{t}^{y}\|_{H}^{2}e^{\gamma t}$
$\leq$ $\|X_{s}^{x}-X_{s}^{y}\|_{H}^{2}e^{\gamma s}+2\int_{s}^{t}e^{\gamma r}\langle A(X_{r}^{x})-A(X_{r}^{y}),X_{r}^{x}-X_{r}^{y}\rangle\,\mathrm{d}r+\gamma\int_{s}^{t}e^{\gamma r}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r$
$\leq$ $\|X_{s}^{x}-X_{s}^{y}\|_{H}^{2}e^{\gamma s}+C_{2}\int_{s}^{t}e^{\gamma r}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r+(\gamma-\delta_{2})\int_{s}^{t}e^{\gamma r}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r.$
For any pair of stopping times $0\leq\lambda\leq\Lambda\leq t$, we get that $\mathbb{P}$-a.s.
$\sup_{\lambda\leq\tau\leq\Lambda}\|X_{\tau}^{x}-X_{\tau}^{y}\|_{H}^{2}+(\delta_{2}-\gamma)\int_{\lambda}^{\Lambda}e^{\gamma(r-\Lambda)}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r$
$\leq$ $\|X_{\lambda}^{x}-X_{\lambda}^{y}\|_{H}^{2}+\sup_{\lambda\leq\tau\leq\Lambda}\|X_{\tau}^{x}-X_{\tau}^{y}\|_{H}^{2}\int_{\lambda}^{\Lambda}C_{2}e^{\gamma(r-\Lambda)}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r.$
After taking the expected value, we get by the stochastic Gronwall’s Lemma [3, Lemma A.1], for $\kappa:=\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}$, and for any $R>0$, by Proposition 5.3,
($\mathbb{P}\left(\left\{\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H}\geq\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}\right\}\right)$
$\leq$ ($\mathbb{P}\left(\left\{\sup_{0\leq r\leq t}\|X_{r}^{x}-X_{r}^{y}\|^{2}_{H}+(\delta_{2}-\gamma)\int_{0}^{t}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r\geq\kappa\right\}\right)$
@@
$\leq$ $\frac{4}{\kappa}e^{4R}\|x-y\|_{H}^{2}+\mathbb{P}\left(\left\{\int_{0}^{t}C_{2}e^{\gamma(r-t)}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r\geq R\right\}\right)$
$\leq$ $\frac{4}{\kappa}e^{4R}\|x-y\|_{H}^{2}+\frac{1}{R}\mathbb{E}\left[\int_{0}^{t}C_{2}e^{\gamma(r-t)}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r\right]$
$\leq$ $\frac{36\|F\|^{2}_{\operatorname{Lip}}}{\varepsilon^{2}}e^{4R}\delta^{2}+\frac{2C_{2}}{\lambda_{0}\delta_{1}(\beta+2)R}\left(\|x\|_{H}^{\beta+2}+\frac{\mathfrak{c}}{\gamma}\right).$
First, choose $R>0$ such that the second term is less than $\frac{\varepsilon}{6\|F\|_{\infty}}$. Then, choose $\delta>0$ such that the first term is less than $\frac{\varepsilon}{6\|F\|_{\infty}}$. By (5.20), we get that,
$|P_{t}F(x)-P_{t}F(y)|<\frac{\varepsilon}{3}+2\|F\|_{\infty}\frac{2\varepsilon}{6\|F\|_{\infty}}=\varepsilon,$
whenever $\|x-y\|_{H}<\delta$. As $\delta>0$ can be chosen independent of $t\geq 0$, the claim is proved. ∎
@@
Assume that Hypotheses (A)(B)(C)(D)(E), and (F) are valid. Then the semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) is a stochastically continuous Markovian Feller semigroup.
@@
$\leq$ $2\sup_{y\in H}\|F_{m}(y)-F(y)\|_{H}+\frac{\varepsilon}{3}=\varepsilon.$
Let us prove the stochastic continuity. Let $x\in H$. By [20, Proposition 2.1.1], it is sufficient to prove that
@@
$\mathbb{E}\left[\|X_{t}^{x}-x\|^{\beta+2}_{H}\right]\leq\mathfrak{c}t\longrightarrow 0,$
@@
The Markov property follows by the same arguments as in [56, Proposition 4.3.5], see also [34, Section 6.4]. ∎

### v4

@@ 5.1. Main estimate
$\exp\left(\|X_{t\wedge\tau_{R}}^{(n)}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{s}^{t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r\right)$
$\leq$ $\exp\left(\|X_{s}^{(n)}\|_{H}^{\beta+2}+((t\wedge\tau_{R})-s)(c_{1}+c_{2}+c_{3})\right)$
$\times\exp\left((\beta+2)\int_{s}^{t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle\right$.
$\left.-\frac{(\beta+2)^{2}}{2}\int_{s}^{t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|Q_{n}B\tilde{Q}_{n}\|_{L_{2}(U,H)}^{2}\,\mathrm{d}r\right)$
$\leq$ $\exp\left(\|X_{s}^{(n)}\|_{H}^{\beta+2}+(t-s)(c_{1}+c_{2}+c_{3})\right)$
$\times\exp\left((\beta+2)\int_{s}^{t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle\right$.
$\left.-\frac{(\beta+2)^{2}}{2}\int_{s}^{t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|Q_{n}B\tilde{Q}_{n}\|_{L_{2}(U,H)}^{2}\,\mathrm{d}r\right).$
The stopped exponential is an exponential martingale which has expectation less or equal to one, see e.g. [90]. As it is independent from $X_{s}^{(n)}$, we obtain,
$\mathbb{E}\left[\exp\left(\|X_{t\wedge\tau_{R}}^{(n)}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{s}^{t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r\right)\right]$
$\leq$ $\mathbb{E}\left[\exp\left(\|X_{s}^{(n)}\|_{H}^{\beta+2}+(t-s)(c_{1}+c_{2}+c_{3})\right)\right].$
@@
$\mathbb{E}\left[\exp\left(\|X_{t}^{(n)}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r\right)\right]$
$\leq$ $\mathbb{E}\left[\exp\left(\|X_{s}^{(n)}\|_{H}^{\beta+2}+(t-s)(c_{1}+c_{2}+c_{3})\right)\right].$
In the locally monotone case, by [70, Proof of Theorem 5.1.3] or [16, Lemma 4.4], there exists a subsequence $\{n_{k}\}$ such that $Q_{n_{k}}X^{(n_{k}),x_{n_{k}}}\rightharpoonup^{\ast}X^{x}$ weakly in $L^{\infty}([0,T];L^{2}(\Omega;H))$ and weakly in $L^{\alpha}([0,T]\times\Omega,{\mathcal{B}}{\mathcal{F}},\mathrm{d}t\otimes\mathbb{P};V)$ as $k\to\infty$, where
${\mathcal{B}}{\mathcal{F}}:=\{A\subset[0,T]\times\Omega\,\colon\,\forall t\in[0,T],A\cap([0,t]\times\Omega)\in{\mathcal{B}}([0,t])\otimes{\mathcal{F}}_{t}\}$
[…]
$\mathbb{E}\left[\exp\left(\|X_{t}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\right)\right]$
$\leq$ $\mathbb{E}\left[\exp\left(\|X_{s}\|_{H}^{\beta+2}+(t-s)(c_{1}+c_{2}+c_{3})\right)\right].$
for $\mathrm{d}t$-almost every $0\leq s\leq t\leq T$ in the locally monotone case, and and similarly with $\mathbb{E}$ replaced by $\tilde{\mathbb{E}}$ in the fully locally monotone case. By continuity of the solutions, see [70, 83], the right-hand side depends continuously on $s$ and $t$, we get the result for every $0\leq s\leq t\leq T$ by a simple approximation argument. The second part of the assertion follows easily, by setting $s=0$. The proof is complete. ∎
A fortiori, when examining the proof of Proposition 5.1, we obtain the following. Note that stopping and taking the expectation removes the stochastic integral.
@@
Assume that Hypotheses (A)(B)(C)(D)(F). Let $T>0$ be fixed. Let $\alpha\geq 2$, $\alpha-2\geq\beta\geq 0$ be as in Hypothesis (C). Then for any initial datum $X_{0}=x$ such that $x\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, and every $0\leq s\leq t\leq T$, the solution $(X_{t})_{t\in[0,T]}$ to (2.1) satisfies
(5.16) $\begin{split}&\mathbb{E}\left[\|X_{t}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}\|_{H}^{\beta+2}\right]+(t-s)(c_{1}+c_{2}),\end{split}$
for every $0\leq s\leq t\leq T$, where $\lambda_{i}$, $i=0,1,2,3$, $c_{i}$, $i=1,2$, are as in Hypothesis (H), and $\delta_{1}>0$ is as in Hypothesis (B). Furthermore,
(5.17) $\begin{split}&\mathbb{E}\left[\|X_{t}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{0}^{t}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|x\|_{H}^{\beta+2}\right]+t(c_{1}+c_{2}),\end{split}$
for every $0\leq t\leq T$.
@@ 5.2. Properties of the semigroup
Assume that Hypotheses (A)(B)(C)(D)(E)(F), and (H) are valid. Let $\varepsilon>0$, $F\in\operatorname{Lip}_{b}(H)$, $x\in H$. Then there exists $\delta>0$ such that for any $y\in H$ with $\|x-y\|<\delta$, and any $t\geq 0$, we have that
@@
Note that if $C_{2}=0$, the $e$-property follows as in [42, Equation (5.4)]. Let $\varepsilon>0$, $x,y\in H$, $\|x-y\|<\delta$, $F\in\operatorname{Lip}_{b}(H)$, $t\geq 0$. Let $(X^{x}_{t})_{t\geq 0}$, $(X^{y}_{t})_{t\geq 0}$ denote the solutions to (2.1) with $X_{0}^{x}=x$, and $X^{y}_{0}=y$, respectively.
Note that due to the additivity of the noise, the synchronous coupling (a.k.a. parallel coupling) yields that the map $t\mapsto(X^{x}_{t}-X^{y}_{t})$ is absolutely continuous in $H$. The chain rule [89, Proposition III.1.2] with the help of Hypothesis (C) yields that $\mathbb{P}$-a.s.,
$\frac{\mathrm{d}}{\mathrm{d}t}\|X^{x}_{t}-X^{y}_{t}\|_{H}^{2}=$ $2\langle A(X^{x}_{t})-A(X^{y}_{t}),X^{x}_{t}-X^{y}_{t}\rangle$
$\leq$ $\left(-\delta_{2}+\eta(X^{x}_{t})+\rho(X^{y}_{t})\right)\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H},$
for all $t\in[0,T]$. Similarly, we get (by interchanging the roles of $x$ and $y$) that
$\frac{\mathrm{d}}{\mathrm{d}t}\|X^{y}_{t}-X^{x}_{t}\|_{H}^{2}=$ $2\langle A(X^{y}_{t})-A(X^{x}_{t}),X^{y}_{t}-X^{x}_{t}\rangle$
$\leq$ $\left(-\delta_{2}+\eta(X^{y}_{t})+\rho(X^{x}_{t})\right)\|X_{t}^{y}-X_{t}^{x}\|^{2}_{H},$
for all $t\in[0,T]$. The preceding two differential inequalities with the help of Hypothesis (C) yields
$\frac{\mathrm{d}}{\mathrm{d}t}\|X^{x}_{t}-X^{y}_{t}\|_{H}^{2}\leq$ $\left(-\delta_{2}+\frac{1}{2}C_{2}\|X^{y}_{t}\|^{\alpha}_{V}\|X^{y}_{t}\|^{\beta}_{H}+\frac{1}{2}C_{2}\|X^{x}_{t}\|^{\alpha}_{V}\|X^{x}_{t}\|^{\beta}_{H}\right)$
$\quad\times\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H},$
for all $t\in[0,T]$.
An application of the differential form of Gronwall’s lemma yields
$\|X^{x}_{t}-X^{y}_{t}\|_{H}^{2}\leq$ $\|x-y\|^{2}_{H}\exp\left(-\delta_{2}t+\frac{1}{2}C_{2}\int_{0}^{t}\left[\|X^{x}_{t}\|^{\alpha}_{V}\|X^{x}_{t}\|^{\beta}_{H}+\|X^{y}_{r}\|^{\alpha}_{V}\|X^{y}_{r}\|^{\beta}_{H}\right]\,\mathrm{d}r\right),$
for all $t\in[0,T]$.
Now, by the Cauchy-Schwarz inequality, Proposition 5.1 and Hypothesis (H),
$\mathbb{E}\left[\|X^{x}_{t}-X^{y}_{t}\|_{H}^{2}\right]$
$\leq$ $\|x-y\|^{2}_{H}\exp\left(-\delta_{2}t\right)$
@@
$\times\left(\mathbb{E}\left[\exp\left(C_{2}\int_{0}^{t}\|X^{x}_{t}\|^{\alpha}_{V}\|X^{x}_{t}\|^{\beta}_{H}\,\mathrm{d}r\right)\right]\right)^{\frac{1}{2}}\left(\mathbb{E}\left[\exp\left(C_{2}\int_{0}^{t}\|X^{y}_{r}\|^{\alpha}_{V}\|X^{y}_{r}\|^{\beta}_{H}\,\mathrm{d}r\right)\right]\right)^{\frac{1}{2}}$
$\leq$ $\|x-y\|^{2}_{H}\exp\left(-\delta_{2}t+\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\left(\|x\|_{H}^{\beta+2}+\|y\|_{H}^{\beta+2}+2t(c_{1}+c_{2}+c_{3})\right)\right)$
$\leq$ $\|x-y\|^{2}_{H}\exp\left(\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\left(\|x\|_{H}^{\beta+2}+\|y\|_{H}^{\beta+2}\right)\right)$
$\leq$ $\|x-y\|^{2}_{H}\exp\left(\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\left(\|x\|_{H}^{\beta+2}+2^{\beta+1}\|y-x\|_{H}^{\beta+2}+2^{\beta+1}\|x\|_{H}^{\beta+2}\right)\right)$
$<$ $\delta^{2}\exp\left(\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\left(\left(1+2^{\beta+1}\right)\|x\|_{H}^{\beta+2}+2^{\beta+1}\delta^{\beta+2}\right)\right)$
and thus,
$|P_{t}F(x)-P_{t}F(y)|^{2}$
$=$ $|\mathbb{E}\left[F(X^{x}_{t})\right]-\mathbb{E}\left[F(X^{y}_{t})\right]|^{2}$
$\leq$ $\mathbb{E}\left[|F(X^{x}_{t})-F(X^{y}_{t})|^{2}\right]$
$\leq$ $\operatorname{Lip}(F)^{2}\mathbb{E}[\left\|X^{x}_{t}-X^{y}_{t}\|^{2}_{H}\right]$
$<$ $\operatorname{Lip}(F)^{2}\delta^{2}\exp\left(\frac{C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\left(\left(1+2^{\beta+1}\right)\|x\|_{H}^{\beta+2}+2^{\beta+1}\delta^{\beta+2}\right)\right)$
$<$ $\varepsilon^{2},$
when $\delta>0$ is small enough. ∎
@@
Assume that Hypotheses (A)(B)(C)(D)(E)(F), and (H) are valid. Then the semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) is a stochastically continuous Markovian Feller semigroup.
@@
$\leq$ $2\sup_{y\in H}\|F_{m}(y)-F(y)\|_{H}+\frac{\varepsilon}{3}<\varepsilon.$
Let us prove the stochastic continuity. Let $x\in H$. By [24, Proposition 2.1.1], it is sufficient to prove that
@@
$\mathbb{E}\left[\|X_{t}^{x}-x\|^{\beta+2}_{H}\right]\leq(c_{1}+c_{2})t\longrightarrow 0,$
@@
The Markov property follows by the same arguments as in [70, Proposition 4.3.5], see also [42, Section 6.4]. ∎